# DG-DETR: Toward Domain Generalized Detection Transformer

### v1

@@ 1 Introduction
To alleviate this problem, the prominent research direction is unsupervised domain adaptation (UDA) [8, 9, 10, 11, 12, 13], which aims to align the data distribution of a labeled source domain to that of an unlabeled target domain. Although UDA methods have demonstrated promising results, they heavily rely on the assumption that target domain data is available during the adaptation process. Moreover, collecting sufficient training data for all target domains of interest is often impractical [14] due to the cost and effort required, even without annotations.
Domain generalization (DG) [15, 16, 17, 14] has emerged as a feasible solution, aiming to train a model that generalizes well to unseen target domains by learning from multiple observed source domains during training. Most DG methods generally strive to learn domain-invariant feature representations across these source domains, but their performance is sensitive to the diversity and quantity of the available source domains [18, 19], which are often costly and labor-intensive to collect. As a more practical approach, recent research has begun focusing on single-domain generalization (S-DG) [20, 21], which aims to train robust models using only one source domain. However, achieving OOD robustness in this setting remains an open challenge.
@@
Figure 1: Overview of the proposed DG-DETR framework.
@@
    To the best of our knowledge, this is the first study to explore the potential of DETRs for domain generalized object detection.
@@
    We introduce DG-DETR, a simple yet effective plug-and-play method that improves the generalization capability of DETRs.
@@ 2.1 Overview
Figure 2: Visualization of wavelet decomposition. The image is decomposed into low- and high-frequency components, where low-frequency bands capture overall style and high-frequency bands preserve object shapes.
@@ 2.2 Wavelet-guided Style Augmentation Module
To this end, we propose to utilize wavelet transform to disentangle high- and low-frequency features of images. Specifically, given an intermediate CNN feature map $F\in\mathbb{R}^{H\times W\times C}$ with spatial dimensions $H\times W$ and $C$ channels, discrete wavelet transform (DWT) applies four kernels, $LL^{T},LH^{T},HL^{T}$, and $HH^{T}$, where the low-pass filter is defined as $L=\left[\begin{array}[]{ll}1/\sqrt{2}&1/\sqrt{2}\end{array}\right]$ and high pass filter is $H=\left[\begin{array}[]{ll}-1/\sqrt{2}&1/\sqrt{2}\end{array}\right]$. These kernels slide across the input $F$ to decompose it into the four wavelet sub-bands: $F_{ll},F_{lh},F_{hl},$ and $F_{hh}$, each with reduced resolution. In this letter, we denote the low-frequency components $F_{ll}$ as $F_{low}$ and the high-frequency components $[F_{lh},F_{hl},F_{hh}]$ as $F_{high}$, respectively.
@@ 2.3 Domain-agnostic Query Selection
[…]
@@
Table 1: Quantitative comparison of domain generalization performance on the Diverse Weather Dataset (DWD). The results represent mAP (%) for different weather conditions when trained on Daytime-Sunny and tested on Night-Sunny, Dusk-Rainy, Night-Rainy, and Daytime-Foggy.
Night-Sunny Dusk-Rainy Method bus bike car motor person rider truck mAP Method bus bike car motor person rider truck mAP RT-DETR [33] 45.4 41.2 70.5 19.1 48.9 30.8 47.6 43.3 RT-DETR [33] 46.1 23.7 74.0 16.1 36.7 16.3 49.7 37.5 SW [39] 34.6 23.7 65.7 10.2 37.1 19.7 35.2 32.3 SW [39] 40.0 14.2 72.6 6.8 34.7 10.5 41.6 31.5 IBN-Net [35] 36.8 24.0 65.5 13.1 39.9 21.2 38.9 34.2 IBN-Net [35] 43.2 16.0 72.3 14.4 35.4 12.2 44.9 34.1 IterNorm [40] 36.8 31.9 68.4 14.6 44.9 25.7 38.8 37.3 IterNorm [40] 35.4 17.4 69.1 15.7 29.7 11.9 39.7 31.3 NP [23] 49.0 40.5 71.9 20.8 49.2 33.4 50.2 45.0 NP [23] 51.8 28.6 77.4 19.0 42.2 22.3 54.1 42.2 DG-DETR (ours) 51.2 44.7 73.4 23.0 51.9 36.1 52.5 47.6 DG-DETR (ours) 51.0 28.1 77.7 20.1 42.0 19.9 55.7 42.1
Night-Rainy Daytime-Foggy Method bus bike car motor person rider truck mAP Method bus bike car motor person rider truck mAP RT-DETR [33] 27.7 7.8 42.0 0.3 12.9 8.7 24.8 17.7 RT-DETR [33] 32.7 27.5 61.8 30.6 38.1 36.9 25.0 36.1 SW [39] 20.3 6.5 42.9 0.1 10.8 4.1 18.7 14.8 SW [39] 23.5 21.4 55.7 20.4 31.1 32.7 12.9 28.2 IBN-Net [35] 29.3 6.6 42.6 0.9 13.7 8.0 24.9 18.0 IBN-Net [35] 19.9 17.8 47.2 19.3 25.7 24.1 13.0 23.9 IterNorm [40] 18.6 7.4 33.6 0.6 9.1 7.9 12.8 12.9 IterNorm [40] 25.9 25.5 59.2 24.7 37.4 36.1 15.6 32.1 NP [23] 39.9 16.8 50.4 1.0 17.7 8.9 32.0 23.9 NP [23] 33.2 27.1 61.7 29.2 38.3 37.6 25.9 36.2 DG-DETR (ours) 39.1 19.6 51.3 4.5 18.3 14.0 32.7 25.6 DG-DETR (ours) 36.4 28.9 64.4 31.2 40.3 39.9 28.8 38.5
Table 2: In-domain performance comparison on the Daytime-Sunny of the Diverse Weather Dataset (DWD). All models are trained and tested under the same weather condition.
@@
Daytime-Sunny Method bus bike car motor person rider truck mAP RT-DETR [33] 61.3 48.3 82.2 51.7 59.8 49.2 63.1 59.4 SW [39] 55.4 40.4 80.5 42.9 54.5 44.5 55.9 53.4 IBN-Net [35] 56.0 37.3 80.0 42.7 53.9 43.1 57.3 52.9 IterNorm [40] 60.4 47.3 82.4 46.2 59.7 47.9 61.2 57.9 NP [23] 61.1 49.7 82.9 49.6 60.5 51.9 63.4 59.9 DG-DETR (ours) 62.5 50.7 83.5 51.9 61.6 53.3 64.9 61.2
@@
Table 3: Ablation analysis of the proposed components. The results presents the impact of Wavelet-guided Style Augmentation (WaveNP) and Domain-agnostic Query Selection (DAQS) on domain generalization performance.
@@ 3.1 Experimental setup
To evaluate our method, we use the Diverse Weather Dataset (DWD)[22], an urban-scene detection benchmark that includes five different weather conditions: Daytime-Sunny (DS), Night-Sunny (NS), Night-Rainy (NR), Dusk-Rainy (DR), and Daytime-Foggy (DF). DWD collected images from BDD-100k [41], FoggyCityscapes [7], and AdverseWeather [42] datasets. Following [22], we train the model using only the source domain (i.e., Daytime-Sunny) and evaluate it directly on the other adverse weather domains.
In all experiments, we adopt RT-DETR [33] as the base detector. To the best of our knowledge, no prior study has explored domain generalization (DG) in DETR-based object detection. Therefore, we extend existing DG methods [39, 35, 40, 18, 23], originally implemented for CNNs, to RT-DETR. All these methods improve the generalization ability of models by applying feature normalization  [39, 35, 40, 18] or perturbation [23] to the CNN feature extractor. Thus, we directly integrate them into DETR’s feature extractor (i.e., CNN backbone). To evaluate domain generalization performance, we follow [22] and use the Mean Average Precision (mAP) metric. Specifically, we report mAP at an Intersection over Union (IoU) threshold of 0.5 (mAP@0.5). The projection scaling factor (Eq. 6) is set to 1.0.
Figure 5: Detection results on Night-Sunny, Dusk-Rainy, Night-Rainy, Daytime-Foggy. First Row: RT-DETR [33] results. Second Row: The results with our method.
@@ 3.2 Performance Analysis of Domain Generalization
We also conducted comparative experiments in in-domain scenarios, and the results are reported in Table 2. The results show that DG-DETR significantly outperforms existing methods, demonstrating that the proposed approach improves performance even when the training and test sets come from the same domain.
@@ 3.3 Ablation Analysis
To further evaluate the effectiveness of our key modules and strategies in DG-DETR, we conduct a series of ablation experiments. Table 3 presents the ablation study on each component. As observed, removing any single component degrades DG performance, demonstrating the contribution of each module. We note that Domain-agnostic Query Selection (DAQS) improves model performance even without augmentation or normalization techniques by removing domain-induced biases from object queries.

### v2

@@ 1 Introduction
To alleviate this problem, the prominent research direction is unsupervised domain adaptation (UDA) [8, 9, 10, 11], which aims to align the data distribution of a labeled source domain to that of an unlabeled target domain. Although UDA methods have demonstrated promising results, they heavily rely on the assumption that target domain data is available during the adaptation process. Moreover, collecting sufficient training data for all target domains of interest is often impractical [12] due to the cost and effort required, even without annotations.
Domain generalization (DG) [13, 14, 15, 12] has emerged as a feasible solution, aiming to train a model that generalizes well to unseen target domains by learning from multiple observed source domains during training. Most DG methods generally strive to learn domain-invariant feature representations across these source domains, but their performance is sensitive to the diversity and quantity of the available source domains [16, 17], which are often costly and labor-intensive to collect. As a more practical approach, recent research has begun focusing on single-domain generalization (S-DG) [18, 19], which aims to train robust models using only one source domain. However, achieving out-of-distribution (OOD) robustness in this setting remains an open challenge.
@@
Figure 1: Overview of the proposed DG-DETR framework. The backbone extracts multi-scale features, represented as $s_{1}$, $s_{2}$, and $s_{3}$, from its last three stages. These features are processed by the encoder and fed into the Domain-agnostic Query Selection (DAQS) module. The DAQS module explicitly removes domain-induced biases from object queries, providing high-quality initial queries for the decoder.
@@
    To the best of our knowledge, this is the first study to explore the potential of DETRs for domain generalized object detection and to propose DG-DETR as a simple yet effective plug-and-play solution.
@@
2 Related Work
2.1 Single-Domain Generalization for Object Detection
To improve model robustness in unseen target domains, the single-domain generalization (S-DG) setting is practical as it addresses the common issue of data scarcity from multiple domains. Recent advances in S-DG can be categorized into four main approaches: data augmentation [22, 31], feature augmentation [29, 30, 32], domain-invariant feature learning [33, 20], and architectural improvements [34]. The most common approach is data augmentation, which aims to diversify the distribution of source data. Lee et al. [22] use object-aware augmentation and contrastive loss to improve model robustness. DivAlign [31] employs carefully selected augmentations to diversify the source domain and further aligns detections across multiple augmented views based on classification and localization. In addition to data augmentation, feature augmentation methods improve generalization by perturbing feature statistics [29], mixing novel styles [30], or leveraging vision-language models for semantic augmentation [32]. Other approaches aim to learn domain-invariant representations or to improve network architectures for enhanced generalization. Wu et al. [20] proposed a cyclic-disentangled self-distillation (CDSD) to extract domain-invariant representations, while G-NAS [34] uses neural architecture search to find more generalizable model structures. While these approaches have advanced single-domain generalization for CNN-based detectors, they are not directly tailored to transformer-based detectors. In this work, we address this gap by introducing DG-DETR with domain-agnostic query selection and wavelet-based feature perturbation to improve OOD robustness in the single-domain generalization setting.
2.2 End-to-end Object Detectors
Transformer-based detectors have recently emerged as a powerful alternative to CNN-based detectors by directly formulating object detection as a set prediction problem. Carion et al. [4] first propose the transformer-based end-to-end detector, DEtection TRansformer (DETR), which eliminates hand-crafted components but suffers from slow convergence. To address this, Deformable-DETR [35] introduces multi-scale deformable attention, significantly improving training efficiency and detection accuracy. Subsequent advances further refine query design, such as dynamic anchor boxes queries [36], denoising queries [37, 38], and adopt hybrid matching strategies [39]. For deployment efficiency, RT-DETR [40] further demonstrates that real-time and end-to-end detection can be achieved through an efficient hybrid encoder and an IoU-aware query selection strategy.
@@ 3.1 Overview
Figure 2: Visualization of wavelet decomposition. The image is decomposed into low- and high-frequency components. In the Discrete Wavelet Transform (DWT) result, $L$ represents the low-frequency approximation, capturing the overall style. LH, HL, and $H$ represent the high-frequency detail coefficients (horizontal, vertical, and diagonal details, respectively), which primarily preserve object shapes and edge information.
@@ 3.2 Wavelet-guided Style Augmentation Module
To this end, we propose to utilize wavelet transform to disentangle high- and low-frequency features of images. Specifically, given an intermediate CNN feature map $F\in\mathbb{R}^{H\times W\times C}$ with spatial dimensions $H\times W$ and $C$ channels, discrete wavelet transform (DWT) applies four kernels, $LL^{T},LH^{T},HL^{T}$, and $HH^{T}$, where the low-pass filter is defined as $L=\left[\begin{array}[]{ll}1/\sqrt{2}&1/\sqrt{2}\end{array}\right]$ and high pass filter is $H=\left[\begin{array}[]{ll}-1/\sqrt{2}&1/\sqrt{2}\end{array}\right]$. These kernels slide across the input $F$ to decompose it into the four wavelet sub-bands: $F_{ll},F_{lh},F_{hl},$ and $F_{hh}$, each with reduced resolution. In this letter, we denote the low-frequency components $F_{ll}$ as $F_{low}$ and the high-frequency components $F_{lh},F_{hl},F_{hh}]$ as $F_{high}$, respectively.
@@ 3.3 Domain-agnostic Query Selection
To address this problem, we propose a Domain-agnostic Query Selection which removes style-induced biases from object latent representations. As illustrated in Fig. 4, our method leverages the orthogonality in the latent space to project queries onto style-irrelevant semantic axes. Given style-related statistics (i.e., $\mu_{s}\text{~and~}\sigma_{s}$), we construct a latent style representation as follows:
@@
Table 1: Quantitative comparison of domain generalization performance on the Diverse Weather Dataset (DWD). The results represent mAP (%) for different weather conditions when trained on Daytime-Sunny and tested on Night-Sunny, Dusk-Rainy, Night-Rainy, and Daytime-Foggy. Bold numbers indicate the best performance, while underlined numbers denote the second-best performance.
Night-Sunny Dusk-Rainy
@@
Method Backbone Detector bus bike car motor person rider truck mAP bus bike car motor person rider truck mAP
Faster R-CNN [1] ResNet-101 - 37.7 30.6 49.5 15.4 31.5 28.6 40.8 33.5 36.8 15.8 50.1 12.8 18.9 12.4 39.5 26.6
RT-DETR [40] ResNet-50 - 45.4 41.2 70.5 19.1 48.9 30.8 47.6 43.3 46.1 23.7 74.0 16.1 36.7 16.3 49.7 37.5
CDSD [20] ResNet-101 Faster-RCNN 40.6 35.1 50.7 19.7 34.7 32.1 43.4 36.6 37.1 19.6 50.9 13.4 19.7 16.3 40.7 28.2
CLIP-Gap [32] ResNet-101 Faster-RCNN 37.7 34.3 58.0 19.2 37.6 28.5 42.9 36.9 37.8 22.8 60.7 16.8 26.8 18.7 42.4 32.3
SW [44] ResNet-50 RT-DETR 34.6 23.7 65.7 10.2 37.1 19.7 35.2 32.3 40.0 14.2 72.6 6.8 34.7 10.5 41.6 31.5
IBN-Net [42] ResNet-50 RT-DETR 36.8 24.0 65.5 13.1 39.9 21.2 38.9 34.2 43.2 16.0 72.3 14.4 35.4 12.2 44.9 34.1
IterNorm [45] ResNet-50 RT-DETR 36.8 31.9 68.4 14.6 44.9 25.7 38.8 37.3 35.4 17.4 69.1 15.7 29.7 11.9 39.7 31.3
NP [21] ResNet-50 RT-DETR 49.0 40.5 71.9 20.8 49.2 33.4 50.2 45.0 51.8 28.6 77.4 19.0 42.2 22.3 54.1 42.2
DG-DETR (ours) ResNet-50 RT-DETR 51.2 44.7 73.4 23.0 51.9 36.1 52.5 47.6 51.0 28.1 77.7 20.1 42.0 19.9 55.7 42.1
Night-Rainy Daytime-Foggy
Method Backbone Detector bus bike car motor person rider truck mAP bus bike car motor person rider truck mAP
Faster R-CNN [1] ResNet-101 - 22.6 11.5 27.7 0.4 10.0 10.5 19.0 14.5 30.7 26.7 49.7 26.2 30.9 35.5 23.2 31.9
RT-DETR [40] ResNet-50 - 27.7 7.8 42.0 0.3 12.9 8.7 24.8 17.7 32.7 27.5 61.8 30.6 38.1 36.9 25.0 36.1
CDSD [20] ResNet-101 Faster-RCNN 24.4 11.6 29.5 9.8 10.5 11.4 19.2 16.6 32.9 28.0 48.8 29.8 32.5 38.2 24.1 33.5
CLIP-Gap [32] ResNet-101 Faster-RCNN 28.6 12.1 36.1 9.2 12.3 9.6 22.9 18.7 36.1 34.3 58.0 33.1 39.0 43.9 25.1 38.5
SW [44] ResNet-50 RT-DETR 20.3 6.5 42.9 0.1 10.8 4.1 18.7 14.8 23.5 21.4 55.7 20.4 31.1 32.7 12.9 28.2
IBN-Net [42] ResNet-50 RT-DETR 29.3 6.6 42.6 0.9 13.7 8.0 24.9 18.0 19.9 17.8 47.2 19.3 25.7 24.1 13.0 23.9
IterNorm [45] ResNet-50 RT-DETR 18.6 7.4 33.6 0.6 9.1 7.9 12.8 12.9 25.9 25.5 59.2 24.7 37.4 36.1 15.6 36.2
NP [21] ResNet-50 RT-DETR 39.9 16.8 50.4 1.0 17.7 8.9 32.0 23.9 33.2 27.1 61.7 29.2 38.3 37.6 25.9 36.2
DG-DETR (ours) ResNet-50 RT-DETR 39.1 19.6 51.3 4.5 18.3 14.0 32.7 25.6 36.4 28.9 64.4 31.2 40.3 39.9 28.8 38.5
Table 2: The performance comparison with state-of-the-art methods on Cityscapes-C. The Avg. means the average performance across 15 corruption types, each evaluated at 5 different severity levels. Bold numbers indicate the best performance, while underlined numbers denote the second-best performance.
Noise Blur Weather Digital
Method Clean Gauss Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixel JPEG Avg.
SupCon [46] 43.2 7.0 9.5 7.4 22.6 20.2 22.3 4.3 5.3 23.0 37.3 38.9 31.6 40.1 24.0 20.1 20.9
FSCE [47] 43.1 7.4 10.2 8.2 23.3 20.3 21.5 4.8 5.6 23.6 37.1 38.0 31.9 40.0 23.2 20.4 21.0
OA-DG [22] 43.4 8.2 10.6 8.4 24.6 20.5 22.3 4.8 6.1 25.0 38.4 39.7 32.8 40.2 23.8 22.0 21.8
PhysAug [48] 42.6 14.3 17.0 11.9 25.6 19.1 25.5 3.9 8.6 21.3 35.3 39.5 27.5 39.1 28.9 19.9 22.6
RT-DETR [40] 50.4 21.9 26.0 19.2 31.6 33.7 31.0 3.7 14.4 20.6 38.1 47.2 29.5 47.0 47.1 39.8 30.1
DG-DETR (ours) 52.2 28.2 33.2 26.4 37.0 40.6 36.0 5.5 24.5 29.2 43.1 50.9 30.9 50.4 50.1 43.1 35.3
Table 3: In-domain performance comparison on the Daytime-Sunny of the Diverse Weather Dataset (DWD). All models are trained and tested under the same weather condition. Bold numbers represent the best performance, while underlined numbers denote the second-best performance.
Daytime-Sunny
Method bus bike car motor person rider truck mAP
Faster R-CNN [1] 63.4 42.9 53.4 49.4 39.8 48.1 60.8 51.1
RT-DETR [40] 61.3 48.3 82.2 51.7 59.8 49.2 63.1 59.4
CDSD [20] 68.8 50.9 53.9 56.2 41.8 52.4 68.7 56.1
CLIP-Gap [32] 55.0 47.8 67.5 46.7 49.4 46.8 54.7 52.6
SW [44] 55.4 40.4 80.5 42.9 54.5 44.5 55.9 53.4
IBN-Net [42] 56.0 37.3 80.0 42.7 53.9 43.1 57.3 52.9
IterNorm [45] 60.4 47.3 82.4 46.2 59.7 47.9 61.2 57.9
NP [21] 61.1 49.7 82.9 49.6 60.5 51.9 63.4 59.9
DG-DETR (ours) 62.5 50.7 83.5 51.9 61.6 53.3 64.9 61.2
@@
Table 4: Ablation analysis of the proposed components. The results presents the impact of Wavelet-guided Style Augmentation (WaveNP) and Domain-agnostic Query Selection (DAQS) on domain generalization performance. DS, NS, DR, NR, and DF stand for daytime-sunny, night-sunny, dusk-rainy, night-rainy, and daytime-foggy, respectively.
@@
Table 6: Ablation analysis of DG-DETR and its compatibility with different DETR-based detectors. The results show the performance of DG-DETR components integrated with RT-DETR and DINO.
Method Backbone Detector DS NS DR NR DF
RT-DETR [40] ResNet-50 - 59.4 43.3 37.5 17.7 36.1
DG-DETR (ours) ResNet-50 RT-DETR 61.2 47.6 42.1 25.6 38.5
DINO [38] ResNet-50 - 57.5 36.1 29.3 10.5 34.6
DG-DETR (ours) ResNet-50 DINO 58.7 38.8 31.4 13.0 35.9
@@ 4.1 Experimental setup
To evaluate our method, we use the Diverse Weather Dataset (DWD) [20] and the Cityscapes-C [49]. DWD is an urban-scene detection benchmark that includes five different weather conditions: Daytime-Sunny (DS), Night-Sunny (NS), Night-Rainy (NR), Dusk-Rainy (DR), and Daytime-Foggy (DF). It was collected from BDD-100k [50], FoggyCityscapes [7], and AdverseWeather [51]. Following [20], we train the model using only the source domain (i.e., Daytime-Sunny) and evaluate it directly on the other adverse weather domains. Cityscapes-C is a robust detection benchmark built upon the Cityscapes dataset [52] and provides 15 corruption types across 5 severity levels. We use the training set of Cityscapes as the source domain, and consider the corrupted versions of the validation set as unseen target domains.
In experiments, we adopt RT-DETR [40] as the base detector. To the best of our knowledge, no prior study has explored domain generalization (DG) in DETR-based object detection. Therefore, we extend existing DG methods [44, 42, 45, 16, 21], originally implemented for CNNs, to RT-DETR. All these methods improve the generalization ability of models by applying feature normalization  [44, 42, 45, 16] or perturbation [21] to the CNN feature extractor. Thus, we directly integrate them into DETR’s feature extractor (i.e., CNN backbone). To evaluate domain generalization performance, we follow [20] and use the Mean Average Precision (mAP) metric. Specifically, we report mAP at an Intersection over Union (IoU) threshold of 0.5 (mAP@0.5). The projection scaling factor (Eq. 6) is set to 1.0.
Figure 5: Detection results on Night-Sunny, Dusk-Rainy, Night-Rainy, Daytime-Foggy. First Row: RT-DETR [40] results. Second Row: The results with our method. Green bounding boxes indicate cars, and red bounding boxes indicate trucks.
@@ 4.2 Performance Analysis of Domain Generalization
To further validate the robustness of our method, we evaluate its performance on the Cityscapes-C dataset, a benchmark for common corruptions. As shown in Table 2, DG-DETR significantly outperforms both the baseline RT-DETR and existing state-of-the-art methods. Our method achieves an average mAP of 35.3%, a substantial improvement over the baseline (30.1%) and the previous best method (22.6%). These results indicate that DG-DETR effectively generalizes to both adverse weather conditions and diverse common corruptions.
We also conducted comparative experiments in in-domain scenarios, and the results are reported in Table 3. The results show that DG-DETR significantly outperforms existing methods, demonstrating that the proposed approach improves performance even when the training and test sets come from the same domain. A similar trend is observed in the Cityscapes-C dataset, where DG-DETR achieves a performance of 52.2% on the clean (source domain) domain, surpassing the baseline RT-DETR’s 50.4%.
@@ 4.3 Ablation Analysis
To further evaluate the effectiveness of our key modules and strategies in DG-DETR, we conduct a series of ablation experiments. Table 4 presents the ablation study on each component. As observed, removing any single component degrades the DG performance, which demonstrates the contribution of each module. We note that Domain-agnostic Query Selection (DAQS) improves model performance even without augmentation or normalization techniques by removing domain-induced biases from object queries.
@@
We further evaluate the plug-and-play capability of DG-DETR by integrating it with different DETR-based detectors [40, 38]. As shown in Table 6, our method consistently improves the generalization performance of both RT-DETR and DINO, which highlights its compatibility and effectiveness across different frameworks.